# DiffPatch: Generating Customizable Adversarial Patches using Diffusion Model

### v1

@@ Abstract
Physical adversarial patches printed on clothing can easily allow individuals to evade person detectors. However, most existing adversarial patch generation methods prioritize attack effectiveness over stealthiness, resulting in patches that are aesthetically unpleasing. Although existing methods using generative adversarial networks or diffusion models can produce more natural-looking patches, they often struggle to balance stealthiness with attack effectiveness and lack flexibility for user customization. To address these challenges, we propose a novel diffusion-based customizable patch generation framework termed DiffPatch, specifically tailored for creating naturalistic and customizable adversarial patches. Our approach enables users to utilize a reference image as the source, rather than starting from random noise, and incorporates masks to craft naturalistic patches of various shapes, not limited to squares. To prevent the original semantics from being lost during the diffusion process, we employ Null-text inversion to map random noise samples to a single input image and generate patches through Incomplete Diffusion Optimization (IDO). Notably, while maintaining a natural appearance, our method achieves a comparable attack performance to state-of-the-art non-naturalistic patches when using similarly sized attacks. Using DiffPatch, we have created a physical adversarial T-shirt dataset, AdvPatch-1K, specifically targeting YOLOv5s. This dataset includes over a thousand images across diverse scenarios, validating the effectiveness of our attack in real-world environments. Moreover, it provides a valuable resource for future research.
@@ 1 Introduction
Figure 1: Example images from our AdvPatch-1K dataset, showcasing the performance of DiffPatch in various scenarios.
With the rapid advancement of computer vision, object detectors are extensively utilized in real-world scenarios, including fields such as autonomous driving [8] and medical diagnostics [33]. Unfortunately, recent works [7, 47, 15, 45] have demonstrated that deep neural networks are vulnerable to maliciously crafted adversarial examples, resulting in decisions that defy common sense. This poses significant threats to safety-critical applications.
In the field of computer vision, adversarial attacks can be categorized into two distinct groups: digital attacks [17, 7] and physical attacks [14, 27, 32]. Different from digital attacks, which involve introducing small perturbations within a digital domain, physical adversarial attacks operate in the tangible world, aiming to mislead visual systems through modifications to the appearance or positioning of physical objects. The execution of physical adversarial attacks can be accomplished through numerous methods, including the application of adversarial patches onto the surface of an object, alterations in environmental lighting conditions, or the introduction of distracting elements.
For physical-world adversarial patches, previous research [43, 48] has primarily focused on enhancing attack performance, which often requires introducing stronger perturbations that significantly deviate patches from the distribution of natural images, making them appear highly unnatural. While most patch generation methods [46, 23] incorporate a total variation loss to smooth the patches, this fundamental issue remains unresolved, rendering the patches easily identifiable by humans. With the rapid development of image generation techniques [16, 20], naturalistic patch generation methods based on deep image generators have gradually emerged [21, 30]. The patches generated by sate-of-the-art generators possess a more natural appearance, significantly enhancing their stealthiness. However, these methods suffer from limited attack effectiveness compared to unnatural patches. Moreover, these methods have high randomness in the generation process, making it difficult to generate customized patches based on a given image.
@@
In this work, we propose a novel Stable Diffusion [39] based adversarial patch generation method called DiffPatch that is capable of generating a naturalistic and customizable adversarial patch based on a reference image specified by the adversary. Our DiffPatch method works as follows. Firstly, it maps the reference image into the latent space and generates adversarial patches by optimizing the noise latent vectors. It adopts Null-text inversion [34] to achieve a near-perfect reconstruction so as to guarantee that the generated images largely preserve the original semantics. Secondly, it introduces an Incomplete Diffusion Optimization (IDO) strategy to ensure the natural appearance and attack performance of the optimized patches and employs an IoU-Detection loss to accelerate convergence. Furthermore, to generate irregular-shaped patches, it utilizes masks to replace the background and covers the gradients in non-essential areas. This can also help reduce information interference from other regions during the optimization process. Based on DiffPatch, we have created an adversarial patch dataset AdvPatch-1K which consists of over a thousand images, featuring attacks in various physical-world scenarios with customizable patches of diverse styles and shapes. Partial results are shown in Figure 1. Our main contributions are as follows:
@@
  • We propose a novel diffusion-based method called DiffPatch for generating naturalistic adversarial patches based on a reference image. DiffPatch employs Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
@@
  • Experimental results show that our DiffPatch achieves the highest attack success rate (ASR), significantly reducing the mAP of various object detectors, rivaling or even surpassing the current state-of-the-art methods of unnatural adversarial patches.
  • We create physical-world adversarial patches by printing them on T-shirts and collect a dataset, AdvPatch-1K, which consists of 1,131 images with annotations. This dataset covers 9 unique adversarial T-shirt designs across various contexts, including both indoor and outdoor settings as well as individual and group photos.
@@ 2 Related Work
2.1 Physical Adversarial Attack
With the widespread application of deep learning, researchers began to investigate the security of these models in real-world scenarios. This focus led to the development of the Basic Iterative Method (BIM) [27], which generates adversarial samples that retain their adversarial properties even after being printed. To improve the robustness of physical adversarial samples, the Expectation Over Transformation (EOT) approach [2] was introduced. EOT enhances resistance to real-world transformations by simulating various modifications that images may undergo. The Digital-to-Physical (D2P) method [24] further advances real-world attack effectiveness by learning the transformations between digital and physical images. AdvPatch [5] marks a milestone by demonstrating the creation of robust adversarial patches that function reliably across diverse environmental conditions. Following this, AdvYOLO [43] demonstrated an attack on object detectors by applying adversarial patches to clothing surfaces, allowing wearers to evade detection. Several similar patch attack methods have since been proposed [48, 46, 52, 22], posing a significant threat to object detectors. Additionally, other physical adversarial attack methods have employed accessories like glasses [41], lighting [51], and shadows [50].
@@ 2.2 Naturalistic Adversarial Patch
Inspired by the imperceptibility constraints in digital attacks, Universal Physical Camouflage Attacks (UPC) [23] employ the $L_{\infty}$ norm to limit perturbations and maintain the natural appearance of patches. Unlike iterative methods that optimize patches on images, the Naturalistic Adversarial Patch (NAP) [21] utilizes a generative approach. NAP leverages GANs to indirectly optimize the patch by modifying initial latent vectors, resulting in patches with a highly natural appearance. Building on these advancements, several GAN-based generation methods have emerged [13, 28]. Additionally, diffusion model-based methods have gained traction in the vision field [9, 30, 49]. The generative power of diffusion models enables the creation of more naturalistic and effective physical adversarial patches.
@@ 2.3 Diffusion-based Image Editing
Recently, diffusion models have excelled in the field of image generation, with models like DALL-E2 [36], Imagen [40], and Stable Diffusion [39] largely surpassing GANs in performance [12]. Beyond generation tasks, diffusion-based image editing techniques have also become mainstream. Leveraging the cross-attention mechanism, Prompt to Prompt (P2P) [18] introduces an innovative framework that enables image editing solely through text prompts. To address challenges with poor reconstruction during editing, Null-text inversion [34] achieves near-perfect reconstruction by optimizing the unconditional embedding at each denoising step. Text-driven Blended Diffusion [3] employs masks to target specific regions in an image, allowing for background replacement and consistent edits within defined areas. Additionally, some adversarial attack methods [49, 10] leverage image editing techniques to modify images while preserving their natural appearance, creating adversarial examples that are visually coherent yet misleading to detection models.
@@ 3 Proposed Approach
Figure 2 provides an overview of DiffPatch. It optimizes the Null text embeddings along incomplete diffusion trajectories and maps the image to the latent vectors at intermediate timestamps, ensuring that the patch maintains a natural appearance during updates. It then applies the Incomplete Diffusion Optimization (IDO) guided by an IoU-Detection Loss to optimize the latent vectors, resulting in faster convergence and enhanced attack performance. During optimization, we skip the U-Net gradients to avoid memory overflow and utilize target masks to eliminate semantic interference from irrelevant areas, thereby generating more naturalistic adversarial patches. Next, we will introduce the key techniques used in DiffPatch.
@@ 3.1 Image-to-Latent Mapping
Our objective is to achieve a perfect mapping between a given image $\mathcal{I}$ and latent vector $z_{t}$, utilizing the image encoding $z_{0}$ to obtain $z_{t}$, which is essentially the reverse process of DDIM sampling [42]. Based on the assumption that the ODE process can be reversed in the limit of small steps, DDIM inversion can be formulated as:
@@
where $\varepsilon_{\theta}$ is the pre-trained U-Net used to predict noise, $\mathcal{C}$ is the condition embedding of $\mathcal{P}$, and $\alpha_{t}$ is calculated based on the schedule $\beta_{0},\ldots,\beta_{T}\in(0,1)$, i.e., $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i})$. Through the aforementioned formula, we can obtain the latent vector $z_{t}$ and utilize it for image reconstruction, yet this does not guarantee perfect reconstruction.
To enhance the quality of generated outputs, Stable Diffusion incorporates a classifier-free guidance mechanism [19] that concurrently engages in unconditional generation and guided generation based on condition embedding $\mathcal{C}$:
@@
where $w$ represents the guidance scale parameter, and $\phi=\mathcal{E}("")$ is the embedding of a null text. The generated content is highly sensitive to textual prompts. During each step of guided generation, even minor variations in the prompt can lead to deviations in the trajectory. This cumulative error can make reconstruction challenging, and such errors tend to amplify when we perform image editing.
To resolve this issue, we separately optimize the null text embedding $\phi_{t}$ for each timestamp $t$ in the order of the diffusion process $t=T\rightarrow t=1$, using the optimized embedding to guide the process back onto the correct trajectory. Similar to Null-text inversion [34], firstly, we set the guidance scale $w$ to 1 and utilize DDIM inversion to obtain intermediate vectors $z^{*}_{T},\ldots,z^{*}_{0}$ where $z^{*}_{0}=z_{0}$. After that, regarding $z^{*}_{T}$ as the initial noise vector $\bar{z}_{T}$, we optimize the null text embedding $\phi_{t}$ through the application of the following objective function, after each iteration of the DDIM sampling step (adopting the default value of 7.5 for the guidance scale $w$):
@@
where $z_{t-1}$ represents the vector obtained after DDIM sampling, and after each step of optimization is completed, we update $\bar{z}_{t-1}=z_{t-1}$. Finally, we can achieve near-perfect reconstruction by utilizing the initial noise vector $\bar{z}_{T}$ and the optimized unconditional embeddings $\{\phi_{t}\}^{T}_{t=1}$.
@@ 3.2 Incomplete Diffusion Optimization
We propose Incomplete Diffusion Optimization (IDO), which adds adversarial perturbations along the trajectory of image reconstruction. IDO consists of two techniques: 1) IoU-Detection Loss and 2) Incomplete Diffusion Process.
@@
Loss function is crucial in guiding the optimization process. After an intermediate adversarial image is fed into the detector, we obtain the following outputs: 1) $B_{box}$: coordinates of the predicted bounding box; 2) $P_{obj}$: probability indicating the presence of an object within the predicted box; and 3) $P_{cls}$: classification probabilities for various objects. Previous work used Common Detection Loss [43] to guide adversarial patch generation:
@@
where $N$ represents the batch size, and $I^{\prime}_{i}$ refers to the $i$-th image in a batch. For each image $I^{\prime}_{i}$, the detector identifies multiple objects and assigns an index $j$ to each detected object. The images $\{I^{\prime}_{i}\}_{i=1}^{N}$ are obtained by performing operation $\mathcal{T}$ on clean images, which means pasting patches on the surface of the clothes of the person inside.
According to the loss function formulation in Equation (4), the optimization for a given image $I^{\prime}_{i}$ is guided solely by the object with the highest confidence product, aiming to minimize the detection probability for that specific target. However, this strategy introduces several challenges. First, the application of adversarial patches is constrained by the availability and accuracy of dataset labels. Incomplete annotations or misdetections can result in critical surfaces remaining unpatched, undermining the optimization process. Moreover, focusing the loss on a single object per iteration leads to slower convergence and increased instability during training, limiting the overall effectiveness of the approach.
We propose an alternative loss function IoU-Detection Loss to address these issues. Formally, it is defined as:
@@
We found that using the default 50-step sampling to generate patches often leads to a loss of semantic meaning and natural appearance during optimization. Excessive sampling steps can amplify subtle changes in $z_{T}$, causing the image reconstruction trajectory to deviate. During the image-to-latent mapping phase, we can match it with the latent vector $z_{\frac{T}{2}}$ at an intermediate timestamp to mitigate semantic loss. The generation of adversarial patch $z_{p}$ can then be represented as:
@@
where $G$ represents the diffusion model. Generating patches using $z_{\frac{T}{2}}$ shortens the generation path, thereby enhancing the stability of the patches under adversarial guidance. However, as $z_{\frac{T}{2}}$ continues to be updated, it becomes more chaotic, making complete denoising challenging in the later stages of the process. To prevent adversarial perturbations from being noticeable, perturbation values are constrained by an infinity norm in digital attacks. Similarly, we constrain the perturbations $\delta$ applied to the latent vector $z_{\frac{T}{2}}$:
@@
Specifically, during denoising, 25 steps of DDIM sampling are required, where U-Net performs 25 noise predictions. The denoising process involves removing noise at timestamp $t$ using a sampling recurrence formula to obtain the vector $z_{t-1}$, then gradually obtaining $z_{0}$, and finally using an image decoder $\Psi$ to restore the image. When optimizing the vector $z_{T/2}$ with the loss function $\mathcal{L}$, it can be expanded using the chain rule as follows:
$\nabla_{z_{T/2}}\mathcal{L}=\frac{\partial\mathcal{L}}{\partial x}\cdot\frac{\partial x}{\partial z_{p}}\cdot\frac{\partial z_{p}}{\partial z_{0}}\cdot\frac{\partial z_{0}}{\partial z_{1}}\cdot\cdots\cdot\frac{\partial z_{T/2-1}}{\partial z_{T/2}},$ (8)
where $z_{p}$ represents the adversarial patch obtained through the image encoder $\Psi$, i.e., $z_{p}=\Psi(z_{0})$ , and $x$ is the training data with $z_{p}$ attached. For each step of DDIM sampling, it can be expressed as:
@@
Each denoising step requires computing and storing the U-Net gradient, which can lead to memory overflow. However, since the predicted noise $\varepsilon_{\theta}(z_{t},t,\mathcal{C})$ follows a standard normal distribution $\mathcal{N}(0,I)$, we can use the reparameterization trick [10] to bypass the gradient computation of the U-Net and obtain an approximate gradient:

### v3

@@ Abstract
Physical adversarial patches printed on clothing can enable individuals to evade person detectors, but most existing methods prioritize attack effectiveness over stealthiness, resulting in aesthetically unpleasing patches. While generative adversarial networks and diffusion models can produce more natural-looking patches, they often fail to balance stealthiness with attack effectiveness and lack flexibility for user customization. To address these limitations, we propose DiffPatch, a novel diffusion-based framework for generating customizable and naturalistic adversarial patches. Our approach allows users to start from a reference image (rather than random noise) and incorporates masks to create patches of various shapes, not limited to squares. To preserve the original semantics during the diffusion process, we employ Null-text inversion to map random noise samples to a single input image and generate patches through Incomplete Diffusion Optimization (IDO). Our method achieves attack performance comparable to state-of-the-art non-naturalistic patches while maintaining a natural appearance. Using DiffPatch, we construct AdvT-shirt-1K, the first physical adversarial T-shirt dataset comprising over a thousand images captured in diverse scenarios. AdvT-shirt-1K can serve as a useful dataset for training or testing future defense methods.
@@ 1 Introduction
Figure 1: Example images from our AdvT-shirt dataset, showcasing the performance of DiffPatch in various scenarios.
Object detectors powered by deep neural networks (DNNs) are integral to modern computer vision systems, enabling critical applications such as autonomous driving [9] and medical diagnostics [39]. However, DNNs are highly susceptible to adversarial attacks [8, 57, 16, 55], which manipulate their outputs in ways that defy logic or compromise safety. These attacks can cause object detectors to misclassify or fail to detect critical objects. This vulnerability poses significant risks, particularly in safety-critical domains.
Adversarial attacks on vision models can be broadly categorized into digital attacks [18, 8] and physical attacks [15, 30, 37, 32, 54]. Unlike digital attacks, which introduce imperceptible perturbations in the digital domain, physical attacks manipulate real-world objects or environments to deceive vision models. Common physical attack strategies include applying adversarial patches [50], altering lighting conditions [62], or introducing distracting elements [36].
While adversarial patches have shown promise in physical attacks, most existing methods [50, 58] prioritize attack effectiveness at the expense of stealthiness, resulting in patches that appear unnatural and are easily detectable. Although techniques like total variation loss [56, 26] have been employed to smooth patches, they fail to fully address this issue. Recent advances in image generation [17, 23] have enabled the creation of more naturalistic patches [24, 34], but these methods often sacrifice attack effectiveness and lack customization capabilities. For instance, recent work [11] leverages image-latent mapping for adversarial sample generation, but its large perturbations in physical settings lead to poor naturalness or effectiveness.
@@
To address these limitations, we propose DiffPatch, a novel adversarial patch generation framework based on Stable Diffusion [45]. DiffPatch generates naturalistic and customizable adversarial patches from a reference image specified by the adversary. Our approach first maps the reference image into the latent space and optimizes noise latent vectors to create adversarial patches. We employ Null-text inversion [40] to ensure near-perfect reconstruction, preserving the original semantics. Additionally, we introduce an Incomplete Diffusion Optimization (IDO) strategy to balance natural appearance and attack performance, along with an IoU-Detection loss to accelerate convergence. To generate irregular-shaped patches, we use masks to replace the background and suppress gradients in non-essential areas, reducing interference during optimization.
Using DiffPatch, we construct AdvT-shirt-1K, a dataset of over a thousand images featuring customizable adversarial patches in diverse physical-world scenarios. This dataset demonstrates the practicality of our approach and provides a valuable resource for future research.
In summary, our key contributions are as follows:
@@
  • We propose a novel diffusion-based method called DiffPatch for generating naturalistic adversarial patches based on a reference image. DiffPatch integrates Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
@@
  • Experimental results demonstrate that our DiffPatch achieves the highest attack success rate (ASR), significantly reducing the mean average precision (mAP) of various object detectors. Its performance rivals or even surpasses state-of-the-art methods for generating unnatural adversarial patches.
  • We construct AdvT-shirt-1K, a dataset of physical-world adversarial patches printed on T-shirts. It includes 1,131 annotated images featuring 9 unique adversarial designs across diverse scenarios, including indoor and outdoor environments, as well as individual and group photos, providing a valuable resource for future defense research.
@@ 2 Related Work
2.1 Physical Adversarial Attacks
Early research on adversarial attacks primarily focused on digital domains, with methods like Basic Iterative Method (BIM) [30] and Projected Gradient Descent (PGD) [38] generating perturbations that remain effective even after physical transformations. To address real-world challenges, the Expectation Over Transformation (EOT) approach [3] was introduced, simulating environmental variations to enhance robustness. The Digital-to-Physical (D2P) method [27] further improved attack effectiveness by modeling the transition between digital and physical domains. These advancements laid the groundwork for physical adversarial attacks, such as AdvPatch [6], which demonstrated robust adversarial patches effective across diverse conditions. Subsequent work, including AdvYOLO [50], extended these ideas to evade object detectors using clothing-based patches. Other physical attack strategies have leveraged accessories like glasses [47], manipulated lighting [62], and exploited shadows [61], highlighting the expanding scope of real-world adversarial threats.
@@ 2.2 Naturalistic Adversarial Patch
Drawing inspiration from the imperceptibility constraints in digital attacks, Universal Physical Camouflage Attacks (UPC) [26] use the $L_{\infty}$ norm to bound perturbations, ensuring patches retain a natural appearance. Unlike traditional iterative methods that optimize patches directly on images, the Naturalistic Adversarial Patch (NAP) [24] adopts a generative approach. NAP employs GANs to optimize patches indirectly by modifying initial latent vectors, producing patches that are both natural-looking and effective. This approach has spurred the development of several GAN-based methods [14, 31]. More recently, diffusion models have emerged as a powerful tool in the vision domain [10, 34, 59], leveraging their generative capabilities to create even more naturalistic and effective physical adversarial patches. These advancements highlight the growing potential of generative models in crafting stealthy and robust adversarial attacks.
@@ 2.3 Diffusion-based Image Editing
Diffusion models have recently achieved remarkable success in image generation, with models like DALL-E2 [42], Imagen [46], and Stable Diffusion [45] outperforming GANs in various benchmarks [13]. Beyond generation, diffusion-based image editing techniques have gained prominence [48]. For instance, Prompt to Prompt (P2P) [21] leverages the cross-attention mechanism to enable text-driven image editing, allowing precise modifications through textual prompts. To address reconstruction challenges during editing, Null-text inversion [40] achieves near-perfect reconstruction by optimizing unconditional embeddings at each denoising step. Text-driven Blended Diffusion [4] uses masks to target specific regions, enabling background replacement and localized edits while maintaining consistency. These advancements have also inspired adversarial attack methods [59, 11], which utilize image editing techniques to craft visually coherent adversarial examples that deceive detection models while preserving natural appearance.
@@ 3 Proposed Approach
In this section, we begin with an overview of our proposed attack framework, followed by a detailed introduction to its three key components.
@@
Figure 2 illustrates the overall framework of DiffPatch. It optimizes Null-text embeddings along incomplete diffusion trajectories and maps the image to latent vectors at intermediate timestamps, ensuring that the patch retains a natural appearance during updates. It employs Incomplete Diffusion Optimization (IDO) guided by an IoU-Detection Loss to optimize the latent vectors, achieving faster convergence and improved attack performance. To address memory constraints, we skip the U-Net gradients during optimization and use target masks to eliminate semantic interference from irrelevant regions, enabling the generation of more naturalistic adversarial patches. Next, we will introduce the key techniques of DiffPatch.
@@ 3.1 Image-to-Latent Mapping
Our goal is to achieve a precise mapping between a given image $\mathcal{I}$ and its corresponding latent vector $z_{t}$. This is accomplished by leveraging the image encoding $z_{0}$ to obtain $z_{t}$, which is essentially the reverse process of DDIM sampling [49]. Based on the assumption that the ODE process can be reversed in the limit of small steps, DDIM inversion can be formulated as:
@@
where $\varepsilon_{\theta}$ is the pre-trained U-Net for noise prediction, $\mathcal{C}$ is the condition embedding of $\mathcal{P}$, and $\alpha_{t}$ is calculated based on the schedule $\beta_{0},\ldots,\beta_{T}\in(0,1)$, i.e., $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i})$. While this formula allows us to obtain the latent vector $z_{t}$ and reconstruct the image, it does not guarantee perfect reconstruction.
To improve the quality of generated outputs, Stable Diffusion employs a classifier-free guidance mechanism [22] , which combines unconditional generation with guided generation based on condition embedding $\mathcal{C}$:
@@
where $w$ is the guidance scale parameter, and $\phi=\mathcal{E}("")$ is the embedding of a null text. The generated content is highly sensitive to textual prompts, and even minor variations can cause trajectory deviations during guided generation. These deviations accumulate, making reconstruction challenging, especially during image editing.
To address this, we separately optimize the null text embedding $\phi_{t}$ for each timestamp $t$ in the diffusion process $t=T\rightarrow t=1$, guiding the process back to the correct trajectory. Similar to Null-text inversion [40], we first set the guidance scale $w$ to 1 and use DDIM inversion to obtain intermediate vectors $z^{*}_{T},\ldots,z^{*}_{0}$, where $z^{*}_{0}=z_{0}$. Treating $z^{*}_{T}$ as the initial noise vector $\bar{z}_{T}$, we optimize $\phi_{t}$ using the following objective function after each iteration of the DDIM sampling step (with the guidance scale $w$ set to the default value of 7.5):
@@
where $z_{t-1}$ is the vector obtained after DDIM sampling. After each optimization step, we update $\bar{z}_{t-1}=z_{t-1}$. This process enables near-perfect reconstruction using the initial noise vector $\bar{z}_{T}$ and the optimized unconditional embeddings $\{\phi_{t}\}^{T}_{t=1}$.
@@ 3.2 Incomplete Diffusion Optimization
We propose Incomplete Diffusion Optimization (IDO), a method that introduces adversarial perturbations along the trajectory of image reconstruction. IDO comprises two key techniques: 1) IoU-Detection Loss and 2) Incomplete Diffusion Process.
@@
When an intermediate adversarial image is input into the detector, it produces three outputs: 1) $B_{box}$: the coordinates of the predicted bounding box; 2) $P_{obj}$: the probability of an object being present within the predicted box; and 3) $P_{cls}$: the classification probabilities for different objects. Previous work [50] employed the Common Detection Loss to guide adversarial patch generation:
@@
where $N$ is the batch size, and $I^{\prime}_{i}$ denote the $i$-th image in the batch. For each image $I^{\prime}_{i}$, the detector identifies multiple objects, assigning an index $j$ to each detection. The images $\{I^{\prime}_{i}\}_{i=1}^{N}$ are generated by applying an operation $\mathcal{T}$ on clean images, which involves pasting patches onto the clothing of individuals in the images.
The loss function in Eq.(4) optimizes each image $I^{\prime}_{i}$ by focusing solely on the object with the highest confidence product, aiming to minimize its detection probability. However, this approach presents several limitations. First, the effectiveness of adversarial patches depends heavily on the availability and accuracy of dataset labels. Incomplete annotations or misdetections can leave critical surfaces unpatched, hindering the optimization process. Second, concentrating the loss on a single object per iteration slows convergence and increases training instability, reducing the overall effectiveness of the method.
We propose an alternative loss function IoU-Detection Loss to address these limitations. Formally, it is defined as:
@@
We observed that using the default 50-step sampling to generate patches often results in a loss of semantic meaning and a diminished natural appearance during optimization. Excessive sampling steps can amplify subtle changes in $z_{T}$, causing significant deviations in the image reconstruction trajectory. To address this, we employ an incomplete diffusion process, where the image is matched to the latent vector $z_{\frac{T}{2}}$ at an intermediate timestamp during the image-to-latent mapping phase. This approach reduces cumulative errors, preserving the patch’s semantic integrity and enhancing its natural appearance. The generation of the adversarial patch $z_{p}$ can then be represented as:
@@
where $G$ represents the diffusion model. By generating patches using $z_{\frac{T}{2}}$, we shorten the generation path, improving the stability of the patches under adversarial guidance. This method effectively minimizes semantic loss while maintaining naturalness. However, as $z_{\frac{T}{2}}$ is updated, it becomes more chaotic, making complete denoising increasingly challenging in later stages. To prevent noticeable adversarial perturbations, we constrain the perturbations $\delta$ applied to the latent vector $z_{\frac{T}{2}}$:
@@
Specifically, during denoising, 25 steps of DDIM sampling are required, involving 25 noise predictions by the U-Net. The denoising process removes noise at timestamp $t$ using a sampling recurrence formula to obtain the vector $z_{t-1}$, gradually leading $z_{0}$, and finally uses an image decoder $\Psi$ to restore the image. During the gradient update process, each step of DDIM sampling can be expressed as:
@@
Each denoising step requires computing and storing the U-Net gradient, which can lead to memory overflow. To address this, we adopt the approach from previous work [11] by bypassing the gradient computation of the U-Net, thereby obtaining an approximate gradient. When optimizing the vector $z_{T/2}$ with the loss function $\mathcal{L}$, it can be expanded using the chain rule as follows:
@@
where $z_{p}$ represents the adversarial patch obtained through the image encoder $\Psi$, i.e., $z_{p}=\Psi(z_{0})$ , and $x$ is the training data with $z_{p}$ attached.